OpenStudy (anonymous):

f(x)=(x-1)^3(x+3)(x-4)^2. Determine how to find the x and y intercepts and the multiplicity of zeros.

OpenStudy (amistre64):

well, when x=0, what value do we obtain for y? and the zeros refer to the values of x that make any factor equal out to zero ...

OpenStudy (amistre64):

the exponents are shorthand for how many times a factor appears, so by default they also define the number of times (multiplicity) a value of x produces a zero

OpenStudy (anonymous):

I got -48 for y

OpenStudy (amistre64):

(-1)^3 (+3) (-4)^2 -1 3 16 ; -48 for y intercept is good

OpenStudy (amistre64):

now we should know by now that: x - r = 0, when x=r so what would you say are the zeroes?

OpenStudy (anonymous):

Im sorry where are you getting r?

OpenStudy (amistre64):

r is some random general variable expressing any value whatsoever that you can think of as opposed to just the specific 3 values presented in the problem. x - r = 0 , anytime x = r, therefore; x-3 = 0 when x=3 i good trick is simply to swap signs on that y intercept parts from: (-1)^3 (+3) (-4)^2 to: (+1)^3 (-3) (+4)^2 this defines our zeros AND the exponents define the multipicities

OpenStudy (anonymous):

so the zeros would be 1, -3 and 4?

OpenStudy (amistre64):

yes :) and the exponents attached to them are their multiplicities

OpenStudy (amistre64):

1^3 = 1*1*1 , so a multiplicity of 3 -3^1 = -3 , so a multiplicity of 1 4^2 = 4*4 , so a multiplicity of 2

OpenStudy (anonymous):

so I have the y intercept as -48 but the x intercept I would set the equation equal to zero right?

OpenStudy (amistre64):

yes, for x intercepts you would set the equation to zero. and since it is already in a factored form, and you are just mulitplying things together, we should know by now that anything times 0 is equal to 0. which is why its simplicity itself to say that the roots, or zeros, are such that "x = r"

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so it would be 0=(x-1)^3(x+3)(x-4)^2

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

can you help me with the first step in solving that?

OpenStudy (amistre64):

i already have. spose we rewrite it as: 0 = a*b*c ; what would you say would fit this equation?

OpenStudy (anonymous):

0

OpenStudy (amistre64):

if a = 0, it is true since 0 = 0*b*c if b = 0, it is true since 0 = a*0*c if c = 0, it is true since 0 = a*b*0 therefore, we can assess each factor and equate them to zero.

OpenStudy (amistre64):

0 = (x-1)^3 (x+3) (x-4)^2 when x-1 = 0, solve for x when x+3 = 0, solve for x when x-4 = 0, solve for x

OpenStudy (anonymous):

oh so the zeros are the x intercepts?